OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF. The following ratios of corresponding sides are equal 8/12=16/24=(4x+5)/(4x+11)

OpenStudy (anonymous):

x= ?

OpenStudy (anonymous):

Ok! Hello again! So we know that the ratio that we need is 8/12 (or 2/3) right? \[\frac{4x+5}{4x+11}=\frac{2}{3}\]

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

Now we just solve for x. So multiply bit sides by \(4x+11\) gives what?

OpenStudy (anonymous):

8x/3+22/3 thats just a guess im not exaclty sure

OpenStudy (anonymous):

Well multiplying both sides by \(4x+11\) will cancel on the right hand side and will remain on the left. Something like\[4x+5 = \frac{2}{3}(4x+11)\]With me so far?

OpenStudy (anonymous):

ok yes that makes more sense

OpenStudy (anonymous):

Sorry so yes you are right, however it is worth not multiplying out until later.

OpenStudy (anonymous):

Now we can multiply both sides by 3 giving: \[3(4x+5) = 2(4x+11)\]

OpenStudy (anonymous):

Does that make sense so far?

OpenStudy (anonymous):

to get rid of the fraction yes im with you

OpenStudy (anonymous):

Brilliant, now we simply multiply through to remove the brackets and solve for x. \[12x+15 = 8x+22\]\[4x = 7\]\[x = \frac{4}{7}\] Take a moment to make sure you understand these steps

OpenStudy (anonymous):

when i put in 4/7 for the answer it said I was wrong

OpenStudy (anonymous):

Oh sorry, I made a typo\[4x = 7\]\[x = \frac{7}{4}\] Hopefully you can see where I went wrong?

OpenStudy (anonymous):

ok yes divide both sides by 4

OpenStudy (anonymous):

We can check our answer by substituting that back in to our equation at the beginning.\[\frac{4x + 5}{4x +11}\]now sub in \(x = \frac{7}{4}\) giving\[\frac{4 \times \frac{7}{4} + 5}{4 \times \frac{7}{4} + 11}\]\[\frac{7+5}{7+11}=\frac{12}{18} = \frac{2}{3}\]Hurray! we have made sure our answer is right!

OpenStudy (anonymous):

Let me know if you have any problems with that

OpenStudy (anonymous):

ok thank you very much you're always a great help